in the formula r=s+2t ,if s is three times greater then t,how would i rewrite this ? r=3(s+2t)
Answer by sabanasir(22)

no then u will write
r = s + 2t
here s= three times t
r=3*t + 2t
r=t(3+2)
r=5t